OpenStudy (anonymous):

Find the coordinates of the x-intercept. -x - 2y = 6

OpenStudy (anonymous):

Turn the equation into y = mx + c So add 2y to both sides and take 6 from both sides to get -x - 6 = 2y Then divide both sides by 2 to get y = -x/2 - 3. So -3 is the intercept.

OpenStudy (anonymous):

So it is (-3,0)

OpenStudy (anonymous):

(0, -3)

OpenStudy (anonymous):

when finding: x-intercept, set y=0 and solve for x; y-intercept, set x=0 and solve for y.

OpenStudy (anonymous):

since you wanted the x-intercept, let y=0 then solve for x....

OpenStudy (anonymous):

thanks, so if I am trying to find x-intercept for -x + 5y = -15, the answer would be (0,15) just like the other one

OpenStudy (anonymous):

Crap sorry, I told you how to get the y intercept. Forget what I said :P

OpenStudy (anonymous):

It would be (15,0) for that one because it would be -x = -15 so x = 15

OpenStudy (anonymous):

so the answer is (-3,0) if I am looking for the x intercept for the first question

OpenStudy (anonymous):

This one would be (-6,0) because -x=6 and x=-6

OpenStudy (anonymous):

ok. Thank you so much

OpenStudy (anonymous):

Put y equal to zero to find the x intercept as when y = 0 , you are on the x axis.